OpenStudy (anonymous):

draw a graph and find the intercept x-y=7 3x+y=1

OpenStudy (alfie):

I guess these two are different functions. Y intercept means where the function crosses the Y axis. The y axis has x = 0 equation, so you put in a system x-y = 7 x = 0 Getting y = -7 . Do the same stuff for the second one :)

OpenStudy (anonymous):

I'm trying to find where the two lines intercept, not the y intercept

OpenStudy (alfie):

Oh lol, I read Y intercept for some reason :D... Well, then it's even easier, just do a system between these two. x = 7+y 3(7+y) +y = 1 x=7+y 21 + 4y = 1 You can continue by yourself.. :)

OpenStudy (alfie):

x=7+y 4y = -20 x = 2 y= -5

OpenStudy (alfie):

Now, if you want to trace a graphic, you will have to trace two lines. You know how to do that? You take the first line, give it a random x, and then you find your Y. Here's your first point. Do the same stuff again, you'll find another point. Between two points there it goes just one line, and you just connect these points getting your draw :).